OpenStudy (anonymous):

How do I find the equation of this hyperbola? Given: One focus at (0,25) Asymptotes intersecting at (0,5) An asymptote passing through (24,13) Those are the only given information. (: Thank you.

OpenStudy (anonymous):

do you know the standard form for the equation of a hyperbola?

OpenStudy (anonymous):

\[\frac{(x-x_0)^2}{a^2}-\frac{(y-y_0)^2}{b^2}=1\]

OpenStudy (anonymous):

since the focus is given above the center (0, 5), the equation should be where the y^2 term is positive... \(\large \frac{(y-y_0)^2}{a^2}-\frac{(x-x_0)^2}{b^2}=1 \)

OpenStudy (anonymous):

Yes, I know the standard form. :) Thanks! Now I know the intersection of asymptotes is the center. :)

OpenStudy (anonymous):

Hi again, :) I already know that I'll use the equation typed up there ^ But after I substitute 0 and 5 to x0 and y0 in the numerator, what will I do next? :(

OpenStudy (anonymous):

I have already solved it. Thanks for your help! :)